What is the equation of this slope? m=9,b=7
Answer by addingup(3677)

slope-intercept: y = mx+b, where m is the slope and b the intercept. What is intercept? it's the point where the line crosses y (see graph below).
In your problem:
m = 9 and b = 7:
y = 9x+7
In order to graph this equation we have to find two ordered pairs (x,y)(x,y)because you need two points in order to draw a line.
Let's find y, let's set x to 0 (0, y)
y = 9(0)+7 Any number times 0 = 0
y = 0+7 = 7
So our first pair is (0, 7)
Now set x to 1 and get a second y so we have a second ordered pair:
y = 9(1)+7
y = 9+7 = 16
Now we have a second pair: (1, 16)
See the graph: